1.Calculate the Support Moment of the shoulder for a worker holding a 3kg drill at shoulder height. The length of the worker’s arm is .85m. Gravity=9.81m/s2

a.What joint is the axis of rotation?


b.What is the force (N) of the load?


c.What is the distance of the load from the axis of rotation?


d.What is the Moment of force of the load on the body?


e.What is the Support Moment the shoulder must exert to avoid movement?

1
Expert's answer
2022-03-08T12:19:26-0500

"\text{elbow joint,}\\ F=mg=29.4~N,"

"l'=\frac l2=0.43~m,"

"M=F\frac l2=12.5~N\cdot m,"

"M_1=-M=-12.5~N\cdot m ."